OpenStudy (anonymous):

log 2 (y+2)-log 2 (y-2)=1

OpenStudy (anonymous):

(y+2)/(y-2) = 2 y+2 = 2y-4 2 = y-4 y = 6 y=6 log2(6+2)-log2(6-2)=1 log2(8)-log2 (4)=1 3-2=1

OpenStudy (anonymous):

I think that I am correct on this